# SymBoltz.jl: a symbolic-numeric, approximation-free and differentiable linear Einstein-Boltzmann solver

### v1

@@
SymBoltz is a new Julia package that solves the linear Einstein-Boltzmann equations. It features a symbolic-numeric interface for specifying equations, is free of approximation switching schemes and is compatible with automatic differentiation. Cosmological models are built from replaceable physical components in a way that scales well in model space. The modeler should simply write down their equations, and SymBoltz solves them and eliminates much of the friction in the process. SymBoltz enables up to $100\times$ shorter model definitions compared to browsing equivalent files in CLASS. Symbolic knowledge enables powerful automation of tasks, such as separating computational stages like the background and perturbations, generating the Jacobian matrix and its sparsity pattern, and interpolating arbitrary expressions from the solution. Modern implicit solvers integrate the full stiff equations at all times, reducing slowdowns by taking long time steps, reusing the Jacobian and LU-factorizing it over several time steps, and using fast linear system solvers. Automatic differentiation gives exact derivatives of any output with respect to any input, which is important for gradient-based Markov chain Monte Carlo methods in large parameter spaces, training of emulators, Fisher forecasting and sensitivity analysis. These features are useful in their own rights, but also reinforce each other in a synergy. Results agree with established codes like CLASS and CAMB. With more work, SymBoltz can grow into an integrated symbolic-numeric cosmological modeling environment with a large library of models that delivers differentiable output as fast as other codes. SymBoltz is available at https://github.com/hersle/SymBoltz.jl with single-command installation and extensive documentation, and welcomes questions, suggestions and contributions.
@@
On the theoretical side, this drives a search for the true cosmological model through modifications to $\mathrm{\Lambda CDM}$ (Bull et al., 2016). Efficiently exploring the space of models benefits from numerical tools that are easy to modify. Minimizing friction in the modeling process encourages relaxing model-dependent approximations, creating user-friendly interfaces and structuring codes in modular components that are easy to replace. Some of the most important such tools in the cosmological modeling toolbox are linear Einstein-Boltzmann solvers ( ?Boltzmann codes?) like CAMB (Lewis et al., 2000) and CLASS (Lesgourgues, 2011a).
@@
On the observational frontier, next-generation surveys like the Square Kilometer Array (Dewdney et al., 2009), Vera C. Rubin Observatory (LSST Science Collaboration, 2009), Dark Energy Spectroscopic Instrument (DESI Collaboration, 2016), Simon’s Observatory (The Simons Observatory Collaboration, 2019) and Euclid (Euclid Collaboration, 2025) promise even more precise data. Setting models apart with upcoming data involves both theoretical model parameters and experimental nuisance parameters that live side-by-side in large $O(100)$-dimensional spaces (Piras et al., 2024). In high dimensions, modern Markov chain Monte Carlo methods like Hamiltonian Monte Carlo and the No-U-Turn Sampler (Hoffman & Gelman, 2011) beat the traditional Metropolis-Hastings algorithm (Hastings, 1970). These methods explore parameter space more efficiently, but need both the likelihood and its derivatives with respect to parameters. Differentiability is also increasingly important in other applications, such as forward-modeling field-level inference using simulations with initial conditions from Boltzmann solvers (Seljak et al., 2017). Automatic differentiation is a way to compute derivatives that is more accurate and can be faster than finite differences, which is an approximate and brute-force method. A differentiable Boltzmann solver is therefore a cornerstone in the next-generation cosmological modeling toolbox.
@@
SymBoltz is a new Boltzmann solver that aims to fill these gaps, featuring a convenient symbolic-numeric interface, approximation-freeness and differentiability. At its core, a Boltzmann code solves the gravitational (Einstein) equations for a gravitational theory coupled to some particle species described by thermodynamic (Boltzmann) equations up to first perturbative order around a homogeneous and isotropic universe. It can predict cosmic microwave background (CMB), baryon acoustic oscillation (BAO) or supernova (SN) observations; or be used in longer computations, such as to generate initial conditions to non-linear $N$-body simulations of large-scale structure.
@@
This article is structured as follows. Section 1 revisits the historical development of Boltzmann codes, sketches their structure, reviews methods for computing derivatives and motivates SymBoltz. Section 2 maps out the architecture and main features of SymBoltz and how it differs from other codes. Section 3 shows some example usage. Section 4 discusses synergies and tradeoffs in the design. Section 5 concludes with the current state of the code and future potential. The appendices list equations, implementation details, comparison settings and testing methods.
@@ 1 History and motivation
Peebles & Yu (1970) were first to numerically integrate a comprehensive set of linearized Einstein-Boltzmann equations. Their work was refined over several years, and Ma & Bertschinger (1995) established the groundwork for modern Boltzmann solvers with their code COSMICS. Seljak & Zaldarriaga (1996) soon realized that one can integrate photon multipoles by parts to reduce many differential equations to one integral solution and released CMBFAST. Their line-of-sight integration method has become a standard technique that greatly speeds up the calculation, but requires truncating the multipoles in the differential equations. This Fortran codebase has since evolved into CAMB [^1] https://camb.info/ written by Lewis et al. (2000), which is one of the two most used and maintained Boltzmann solvers today. Doran (2005a) ported CMBFAST to C++ with the fork CMBEASY, which structured the code in an object-oriented fashion and improved user-friendliness, but this project is abandoned today.
@@
More recently, the market has seen an influx of alternative solvers that target new numerical techniques, GPU parallelization, differentiability, interactivity and symbolic computation. Refregier et al. (2017) published the first symbolic-numeric Boltzmann code PyCosmo [^3] https://pypi.org/project/PyCosmo/, which automatically generates efficient C++ ODE code from user-provided symbolic equations, performs sparsity optimization on its Jacobian matrix and avoids the use of approximation schemes to speed up and stabilize the integration. Hahn et al. (2024) published the first differentiable Boltzmann solver DISCO-EB [^4] https://github.com/ohahn/DISCO-EB in the JAX framework in Python, and relaxes approximation schemes to avoid complications and overhead when switching ODEs on GPUs. Bolt.jl [^5] https://github.com/xzackli/Bolt.jl by Li et al. (2023) accomplishes much of the same in Julia. SymBoltz.jl [^6] https://github.com/hersle/SymBoltz.jl is inspired by some of these developments.
@@ 1.1 Structure of traditional Boltzmann solvers
In principle, the core task of a Boltzmann solver is thus very simple: to solve these ordinary differential equations (ODEs) ${\mathrm{d}\mathbf{u}}/{\mathrm{d}\tau}=\mathbf{f}(\tau,\mathbf{u};\mathbf{p})$ for some initial conditions $\mathbf{u}(\tau_{0})$ and parameters $\mathbf{p}$, including several perturbation wavenumbers, and compute desired quantities from their solution, like power spectra.
@@
In practice, however, several properties complicate this task. First, the equations separate into computational stages that benefit substantially from being solved sequentially for performance and stability, such as the background and perturbations stages. It is common to solve each stage with interpolated input from the previous stage, and joining these can be cumbersome and fragment code. Second, the set of equations is very long and convoluted. Ideally, parts of the equations should be easily replaceable to accommodate different cosmological (sub)models. Organizing a code with a suitable modular structure that scales well in model space is hard. Third, tension between wildly different timescales that coexist in the equations make them extremely stiff and intractable for standard explicit ODE solvers. This stiffness must be massaged away in the equations or dealt with numerically in other ways. Fourth, the size of differential equations is very large and increases for more accurate description of relativistic species. Typical models with accurate treatment of photons and neutrinos need $O(100)$ equations. Fifth, the perturbations must be solved for many different wavenumbers $k$, and tradeoffs between performance and accuracy must be made.
@@
To overcome these challenges, Boltzmann solvers are typically written in low-level high-performance languages like C, C++ and Fortran. Codes are usually tightly adapted to the pipeline-like computational structure of the problem (e.g. input $\rightarrow$ background $\rightarrow$ thermodynamics $\rightarrow$ perturbations $\rightarrow\ldots\rightarrow$ output in section 3.3 in Lesgourgues (2011a)). This makes sense for programmers and computers, but not always for modelers.
@@
Traditional codes have nonexistent or only very thin abstraction layers around them. To modify them, users must work directly in the low-level numerical interface and should be prepared to get their hands dirty. For example, to implement a new species, one must generally modify the code in many places: input handling for any new parameters, new background equations, new thermodynamics equations, new perturbation equations, joining between each of these stages, output handling, and possibly more. This leads to fragmented code where the changes related to one species are scattered throughout the code, making it challenging and unintuitive to navigate and decipher.
@@
More importantly, this structure scales poorly in model space. Each module becomes more polluted as more components are added. Even if they are deactivated by if statements at runtime, their mere presence in the source code reduces its readability. In turn, the whole code grows into a complex and mysterious beast.
@@
The two-language problem amplifies the problem, as data analysis usually takes place in slower high-level languages like Python. This shapes Boltzmann solvers to rigid pipelines that must compute everything in one shot and avoid interception at all costs, in order to maximize performance in the low-level language before passing the data back to the high-level language. Some features that are really just post-processing of the ODE solutions are appended to the pipeline even if they are peripheral to the core task of an Einstein-Boltzmann solver – which is to solve the Einstein-Boltzmann equations. This includes non-linear boosting and CMB lensing, for example.
These code smells lead to big monolithic ?black boxes? that scale poorly in model space, and whose complexity grows to incorporate features beyond their original core scope. This complexity has even fueled development of specialized AI assistants for CLASS (Casas et al., 2025). While such tools are helpful, we think they are a symptom of unnecessary complexity and poor scaling with the number of models. Of course, there are also some advantages to this structure, as we discuss in section 4.
@@ 1.2 Boltzmann solver approximation schemes
The Einstein-Boltzmann equations are infamously stiff. This is a property of differential equations (3) that means their numerical solution is unstable and requires tiny step sizes with standard explicit Runge-Kutta methods. Stiffness can arise when multiple and very different time scales appear in the same problem. This is very common in cosmology, where particles interact very rapidly in a universe that expands very slowly, particularly in the tightly coupled baryon-photon fluid, for example. Stiff equations are practically impossible to integrate with explicit solvers and require special treatment.
For a long time, Boltzmann solvers have massaged away stiffness with several approximation schemes [^7] Here ?approximations? refers to schemes that switch between different equations at different times. It excludes techniques like multipole truncation and line-of-sight integration, which SymBoltz also uses. in the equations (e.g. Doran (2005b); Blas et al. (2011)):
@@
These usually involve switching from one set of equations to another when some control variable that measures the applicability of the approximation crosses a threshold. This can change the unknowns in the ODE and require reinitializing it. The approximations help stability and sometimes performance of the equations and allow the use of explicit solvers.
However, approximations put much more strain on the modeler to derive and validate versions of their equations in different regimes. Furthermore, this process should generally be repeated when further changes are made to the model. The numerics also become more complicated: timeseries from each ODE solution must be stitched together, tolerances are duplicated for each separate ODE system that may respond differently to them, ODE integrators must be reinitialized, and so on.
It is a misunderstanding in the literature that it is ?difficult? or ?impossible? to solve the Einstein-Boltzmann equations without using approximations, or that they are ?unavoidable? (Ma & Bertschinger, 1995; Lewis, 2025; Lesgourgues, 2011a, respectively). This statement must only be understood in the context of explicit solvers! Implicit integrators can solve stiff equations, but seem largely underutilized by Boltzmann solvers, perhaps because they are scarce and harder to implement. CLASS has an implicit solver, but does not permit disabling all approximations Blas et al. (2011). On the contrary, new life has been breathed into this field recently, as development of modern implicit solvers and techniques like automatic and symbolic differentiation make implicit solvers more feasible and powerful.
@@ 1.3 Differentiation methods
Derivatives are important in computing. Cosmological applications are no exception. For example, algorithms that optimize likelihoods and Markov chain Monte Carlo (MCMC) samplers for Bayesian parameter inference can take advantage of derivatives of the likelihood with respect to each parameter to intelligently step in a direction where the likelihood or probability increases. In machine learning, the same applies when training neural networks emulators for cosmological observables as functions of cosmological parameters by minimizing scalar loss functions. Some cosmologies are parametrized as boundary-value problems with the shooting method, and use nonlinear root solvers like Newton’s method that need Jacobians. Jacobians are also used by implicit ODE solvers to solve for values at the next time step. Fisher forecasting uses the Hessian (double derivative) of the likelihood with respect to parameters to predict how strong parameter constraints that can be placed by data with given errors. Boltzmann solvers typically save computational resources by interpolating spectra computed on coarse grids of $k$ and $l$ to finer grids, and this can be made more precise with knowledge of derivatives with respect to $k$ and $l$. These are all examples where Boltzmann solvers or applications that use them need derivatives.
There are at least four ways to compute derivatives. Manual differentiation is the exact pen-and-paper method with differentiation rules, but is limited to simple expressions and by human error. Symbolic differentiation automates this process with computer algebra systems, but is inherently symbolic and cannot differentiate arbitrary programs nested with control flow through conditional statements and loops that depend on numerical values. Finite differentiation approximates the derivative $f^{\prime}(x)\approx(f(x+\epsilon/2)-f(x-\epsilon/2))\,/\,\epsilon$ with a small $\epsilon>0$ (here using central differences) by simply evaluating the program several times. This can differentiate arbitrary programs, but is approximate, introduces the step size $\epsilon$ as a hyperparameter that must be tuned for both accuracy and stability, and is a brute-force approach that requires $O(n)$ ($2n$ using central differences) evaluations to compute the gradient of an $n$-variate scalar $f$. Automatic differentiation (e.g. Griewank & Walther, 2008) can be understood by viewing any computer program as a (big) composite function
@@
through the Jacobian $(\boldsymbol{J}_{n})_{ij}=\partial f_{n,i}/\partial f_{n-1,j}$ of every elementary operation to accumulate the derivative of the entire program. This is numerically exact, free of precision parameters, usually requires fewer operations than finite differences, but is perhaps less intuitive, harder to implement and needs the source code to interpret the program (1) in the non-standard way (2).
@@
In practice, both modes are implemented using techniques like operator overloading or source code transformation. The former overloads every operation to accept a special number type that propagates both values and derivatives, such as dual numbers for forward-mode (e.g. Revels et al., 2016). The latter transforms the code for $\boldsymbol{f}$ into another code that computes $\boldsymbol{J}$. In any case, automatic differentiation requires access to the source code!
GravityBaryonsPhotonsMasslessneutrinosMassiveneutrinosCold darkmatterCosmo-logicalconstant Symbolic model BackgroundPerturbations Numerical problem BackgroundPerturbations Solution object CompileSolve
Figure 1: SymBoltz represents cosmological models with equations in symbolic form and grouped by physical components like the spacetime metric, gravity and particle species. This is compiled to a numerical problem that splits the model into background and perturbation stages and generates efficient numerical code for ODEs and their Jacobians. The problem is then solved and the result stored in a solution object that gives convenient access to any quantity defined in the symbolic model.

### v2

@@
SymBoltz is a new Julia package that solves the linear Einstein-Boltzmann equations. It features a symbolic-numeric interface for specifying equations, is free of approximation switching schemes and is compatible with automatic differentiation. Cosmological models are built from replaceable physical components in a way that scales well in model space, or alternatively written as one compact system of equations. The modeler should simply write down their equations, and SymBoltz solves them and eliminates friction in the modeling process. Symbolic knowledge enables powerful automation of tasks, such as separating computational stages like the background and perturbations, generating the analytical Jacobian matrix and its sparsity pattern, and interpolating arbitrary variables from the solution. Implicit solvers integrate the full stiff equations at all times without approximations, which greatly simplifies the code. Performance remains as good as in existing approximation-based codes due to high-order implicit methods that take long time steps, fast generated code, optimal handling of the Jacobian and efficient sparse matrix methods. Automatic differentiation gives exact derivatives of any output with respect to any input, which is important for gradient-based Markov chain Monte Carlo methods in large parameter spaces, training of emulators, Fisher forecasting and sensitivity analysis. The main features form a synergy that reinforces the design of the code. Results agree with established codes to $0.1\%$ with standard precision. More work is needed to implement additional features and for fast reverse-mode automatic differentiation of scalar loss functions. SymBoltz is available at https://github.com/hersle/SymBoltz.jl with single-command installation and extensive documentation, and welcomes all contributions.
@@
On the theoretical side, this drives a search for the true cosmological model through modifications to $\mathrm{\Lambda CDM}$ (Bull et al., 2016). Efficiently exploring the space of models benefits from numerical tools that are easy to modify. Minimizing friction in the modeling process encourages relaxing model-dependent approximations, creating user-friendly interfaces and structuring codes with modular components that are easy to replace. Some of the most important such tools in the cosmological modeling toolbox are linear Einstein-Boltzmann solvers ( ?Boltzmann codes?) like CAMB (Lewis et al., 2000) and CLASS (Lesgourgues, 2011a).
@@
On the observational frontier, next-generation surveys like the Square Kilometer Array (Dewdney et al., 2009), Vera C. Rubin Observatory (LSST Science Collaboration, 2009), Dark Energy Spectroscopic Instrument (DESI Collaboration, 2016), Simon’s Observatory (The Simons Observatory Collaboration, 2019) and Euclid (Euclid Collaboration, 2025) promise more precise data. Setting models apart with upcoming data involves both theoretical model parameters and experimental nuisance parameters that coexist in large $O(100)$-dimensional spaces (Piras et al., 2024). In high dimensions, modern Markov chain Monte Carlo methods like Hamiltonian Monte Carlo and the No-U-Turn Sampler (Hoffman and Gelman, 2011) outperform the traditional Metropolis-Hastings algorithm (Hastings, 1970). They explore parameter space more efficiently, but need both the likelihood and its gradient with respect to parameters. Emulators have become popular to get this from existing non-differentiable codes by training differentiable neural networks to reproduce their output (e.g Bolliet et al., 2024; Bonici et al., 2024). Differentiability is also used in forward-modeling field-level inference with simulations and initial conditions from Boltzmann solvers, for example (Seljak et al., 2017). Automatic differentiation can compute derivatives more accurately and faster than approximate and brute-force finite differences. A differentiable Boltzmann solver is a cornerstone in the modern cosmological modeling toolbox.
@@
SymBoltz is a new Boltzmann solver that aims to meet these needs, featuring a convenient symbolic-numeric interface, approximation-freeness and differentiability. At its core, a Boltzmann code solves the Einstein equations for a gravitational theory coupled to some particle species described by Boltzmann equations up to first perturbative order around a homogeneous and isotropic universe. For example, it predicts cosmic microwave background (CMB), baryon acoustic oscillation (BAO) or supernova (SN) observations, or generates initial conditions for non-linear $N$-body simulations of large-scale structure.
@@
This article is structured as follows. Section 1 motivates SymBoltz by reviewing the historical development of Boltzmann codes and methods for computing derivatives. Section 2 describes the structure and main features of SymBoltz and how it differs from other codes. Section 3 shows example usage. Section 4 compares performance to an existing code. Section 5 describes possibilities for future work. Section 6 concludes with the current state of the code. Appendices A, B and C list equations, implementation details, comparison settings and testing methods.
@@ 1 History and motivation
Peebles and Yu (1970) were first to numerically integrate a comprehensive set of linear Einstein-Boltzmann equations. Their work was refined over several years, and Ma and Bertschinger (1995) established the groundwork for modern Boltzmann solvers with their code COSMICS. Seljak and Zaldarriaga (1996) soon realized that one can integrate photon multipoles by parts to reduce many differential equations to one integral solution and released CMBFAST. Their line-of-sight integration method has become a standard technique that greatly speeds up the calculation, but requires truncating the multipoles in the differential equations. This Fortran codebase has since evolved into CAMB [^1] https://camb.info/ written by Lewis et al. (2000), which is one of the two most used and maintained Boltzmann solvers today. Doran (2005a) ported CMBFAST to C++ with the fork CMBEASY, which structured the code in an object-oriented fashion and improved user-friendliness, but this project is abandoned today.
@@
Recently, several alternative solvers have appeared on the market. They target new numerical techniques, lack of approximations, differentiability, GPU parallelization, interactivity and symbolic computation. PyCosmo [^3] https://pypi.org/project/PyCosmo/ by Refregier et al. (2018) in Python was the first symbolic-numeric Boltzmann code that generates fast C++ code from user-provided symbolic equations, optimizes the sparse and analytical Jacobian matrix and avoids approximation schemes. Bolt.jl [^4] https://github.com/xzackli/Bolt.jl by Li et al. (2023) in Julia is also approximation-free, supports forward-mode automatic differentiation and uses similar differential equations solvers as SymBoltz. DISCO-EB [^5] https://github.com/ohahn/DISCO-EB by Hahn et al. (2024) in the JAX framework in Python is also differentiable and relaxes approximation schemes to avoid overhead from switching equations on GPUs. SymBoltz.jl [^6] https://github.com/hersle/SymBoltz.jl is inspired by all these developments.
@@ 1.1 Structure of traditional Boltzmann solvers
In principle, the core task of a Boltzmann solver is thus very simple: to solve these ordinary differential equations (ODEs) ${\mathrm{d}\boldsymbol{u}}/{\mathrm{d}\tau}=\boldsymbol{f}(\tau,\boldsymbol{u},\boldsymbol{p})$ for some initial conditions $\boldsymbol{u}(\tau_{0})$ and parameters $\boldsymbol{p}$, including several perturbation wavenumbers, and compute desired quantities from their solution, like power spectra.
@@
In practice, however, several properties complicate this task. First, the equations separate into computational stages that benefit substantially from being solved sequentially for performance and stability, such as the background and perturbations stages. It is common to solve each stage with interpolated input from the previous stage, and joining these can be cumbersome and fragment code. Second, the set of equations is very long and convoluted. Ideally, parts of the equations should be easily replaceable to accommodate different submodels for gravity and particle species. Organizing a code with a suitable modular structure that scales well in model space is hard. Third, wildly different timescales coexist in the equations and make them extremely stiff and intractable to solve with standard explicit ODE solvers. This stiffness must be massaged away in the equations or dealt with numerically in other ways. Fourth, the size of differential equations is very large and increases for more accurate description of relativistic species. Typical models with accurate treatment of photons and neutrinos need $O(100)$ equations. Fifth, the perturbations must be solved for many different wavenumbers $k$, and tradeoffs between performance and precision must be made.
@@
To overcome these challenges, most Boltzmann solvers are written in low-level high-performance languages like C, C++ and Fortran. They are tightly adapted to the pipeline-like computational structure of the problem (e.g. input $\rightarrow$ background $\rightarrow$ thermodynamics $\rightarrow$ perturbations $\rightarrow\ldots\rightarrow$ output in Lesgourgues (2011a)). This makes sense for programmers, but does not necessarily provide the simplest interface for modelers.
@@
Traditional codes like CAMB and CLASS have nonexistent or thin abstraction layers. To modify them, users must work directly in the low-level numerical code and have a good understanding of its internal structure. For example, to implement a new species, one must often modify the code in many places: input handling for new parameters, new background equations, new thermodynamics equations, new perturbation equations, joining each of these stages, output handling, and so on. This leads to fragmented code where changes related to one species are scattered throughout the code.
@@
This structure scales poorly in model space. As more species or gravitational theories are added, each module becomes intertwined with code from other physical components. Even if deactivated by if statements at runtime, their mere presence in the source code increases its complexity and reduces its readability.
@@
The two-language problem amplifies the problem, as data analysis often happens in slow high-level languages like Python. This shapes Boltzmann solvers to rigid pipelines that must compute everything at once and avoid interception at all costs, in order to maximize performance in the low-level language before passing output back to the high-level language. Some features are really just post-processing of the ODE solutions, but are appended to the pipeline even if peripheral to the core task of an Einstein-Boltzmann solver – solving the Einstein-Boltzmann equations. Features like non-linear boosting and CMB lensing could be made as smaller and interoperating modular packages, but instead become part of ?all-in-one? Boltzmann codes.
These patterns lead to big monolithic Boltzmann solvers. They become increasingly complex as they incorporate more models and features beyond their original core scope. This complexity has even driven development of specialized AI assistants for CLASS (Casas et al., 2025). While existing Boltzmann solvers are impressively well-engineered and such tools are only helpful, we think these are symptoms of unnecessary complexity.
@@ 1.2 Boltzmann solver approximation schemes
The Einstein-Boltzmann equations are infamously stiff (Nadkarni-Ghosh and Refregier, 2017). This property of differential equations means their numerical solution is unstable with standard explicit integrators and requires tiny step sizes, making them hard to solve. Stiffness can arise when multiple and very different time scales appear in the same problem. This is very common in cosmology, where particles interact very rapidly in a universe that expands very slowly, particularly in the tightly coupled baryon-photon fluid, for example. Stiff equations are practically impossible to integrate with explicit solvers and require special treatment.
For a long time, Boltzmann solvers have massaged away stiffness with several approximation schemes [^7] Here ?approximations? refers to schemes that switch between different equations at different times. It excludes techniques like multipole truncation and line-of-sight integration, which SymBoltz also uses. in the equations (e.g. Doran (2005b); Blas et al. (2011); Lewis (2025)):
@@
  • non-cold dark matter fluid approximation (NCDMFA),
@@
They involve switching from one set of equations to another when some control variable measuring the applicability of the approximation crosses a threshold. This can change the variables in the ODE and require reinitializing it. Approximations can help both speed and stability and allow the use of explicit solvers.
However, approximations put much more strain on modelers to derive and validate several versions of the equations in different regimes. This process may have to be repeated when the model is modified. The numerics also become more complicated: timeseries from each ODE solution must be stitched together, each separate ODE system can use different tolerances, ODE integrators must be reinitialized, and so on.
Another way to integrate stiff equations is to use appropriate implicit solvers. PyCosmo first solved the full stiff Einstein-Boltzmann equations with an implicit integrator, followed by Bolt.jl and DISCO-EB. CLASS also has an implicit solver, but does not permit disabling the tight-coupling approximation, for example. Historically, implicit solvers may have been underutilized because they are harder to implement than explicit solvers and have a reputation for being slower, and due to iterative evolution from initial Boltzmann codes that were centered around approximations. However, new life has recently been breathed into the field of implicit solvers, with development of new methods combined with techniques like automatic and symbolic differentiation that make them more feasible and powerful (e.g. Steinebach, 2023; Ekanathan et al., 2025).
@@ 1.3 Differentiation methods
Derivatives are important in computing. Cosmological applications are no exception. For example, algorithms that optimize likelihoods and Markov chain Monte Carlo (MCMC) samplers for Bayesian parameter inference can take advantage of derivatives of the likelihood with respect to each parameter to intelligently step in a direction where the likelihood increases. In machine learning, the same applies when training neural networks emulators for cosmological observables as functions of parameters by minimizing a scalar loss. Some cosmologies are parametrized as boundary-value problems with the shooting method, and use nonlinear root solvers like Newton’s method that need Jacobians. Implicit ODE solvers also use Jacobians to solve for values at the next time step. Fisher forecasting uses the Hessian (double derivative) of the likelihood with respect to parameters to predict how strong parameter constraints that can be placed by data with some uncertainty. Boltzmann solvers save time by interpolating spectra computed on coarse grids of $k$ and $l$ to finer grids, and this can be made more precise with derivatives with respect to $k$ and $l$. These are all examples where (applications of) Boltzmann solvers need derivatives.
There are at least four ways to compute derivatives.
Manual differentiation is human application of differentiation rules, but is limited to simple expressions and by human error.
Symbolic differentiation automates this process with computer algebra systems, but is inherently symbolic and cannot differentiate arbitrary programs nested with control flow like conditional statements and loops that depend on numerical values.
Finite differentiation approximates the derivative $f^{\prime}(x)\approx(f(x+\epsilon/2)-f(x-\epsilon/2))\,/\,\epsilon$ with a small $\epsilon>0$ (here using central differences) by simply evaluating the program several times. This can differentiate arbitrary programs, but is approximate, introduces the step size $\epsilon$ as a hyperparameter that must be tuned for both accuracy and stability, and is a brute-force approach that requires $O(n)$ evaluations ($2n$ using central differences) to compute the gradient of an $n$-variate scalar $f$.
Automatic differentiation can be understood by viewing any computer program as a (big) composite function
@@
through the Jacobian $(\boldsymbol{J}_{n})_{ij}=\partial f_{n,i}/\partial f_{n-1,j}$ of every elementary operation to accumulate the derivative of the entire program. This is numerically exact, free of precision parameters, usually requires fewer operations than finite differences, but is perhaps less intuitive, harder to implement and needs the source code of the program (1) to interpret it in the non-standard way (2).
@@
In practice, both modes are implemented using techniques like operator overloading or source code transformation. The former specializes every function on a particular ?dual number? that computes both the value and derivative of the function (e.g. Revels et al., 2016). The latter analyzes the source code for $\boldsymbol{f}$ and transforms it into another code that computes $\boldsymbol{J}$. In any case, automatic differentiation requires access to the source code!
GravityBaryonsPhotonsMasslessneutrinosMassiveneutrinosCold darkmatterCosmo-logicalconstant Symbolic model Background: $\boldsymbol{f}$, $\boldsymbol{J}$Perturbations: $\boldsymbol{f}$, $\boldsymbol{J}$ Numerical problem Background: $S(\tau)$Perturbations: $S(\tau,k)$ Solution object CompileSolve
Figure 1: SymBoltz represents cosmological models with symbolic equations grouped in physical components for the metric, gravity and particle species. This is compiled to a numerical problem that splits equations into background and perturbation stages and generates fast code for ODE functions $\boldsymbol{f}$ and Jacobians $\boldsymbol{J}$. The problem is then solved and the result stored in a solution object that gives access to any variable $S$ in the model.